What is the density in g/L of a 14.1 g sample of carbon dioxide gas at 27.2 °C and 3.18 atm?
Answer #1

P = 3.18atm

T= 27.2 oC

= (27.2+273) K

= 300.2 K

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

3.18 atm *44.01 g/mol = density * 0.08206 atm.L/mol.K *300.2 K

density = 5.6812 g/L

Answer: 5.68 g/L